OpenStudy (anonymous):

A rectangle has a perimeter of 26 feet. Twice its length (l) is one foot more than three times its width (w). What is the length of the rectangle, in feet?

OpenStudy (nurali):

Let L represent the length of the rectangle Let W represent the width of the rectangle Perimeter formula: 2L + 2W = 26 Given: 2L = 3W + 1 (I guess this was the hard part) So, 3W + 1 + 2W = 26 \\ solve for W = 5 Therefore 2L + 10 = 26 \\ solve for L 2L=26-10 2L=16 L=8